if you were in a car going 275mph and you slammed on the brakes and it took 3 seconds to get to 83 mph whats the acceleration?
Answer by Alan3354(69443)

if you were in a car going 275mph and you slammed on the brakes and it took 3 seconds to get to 83 mph whats the acceleration?
------------
275 mi/hr = 275*22/15 ft/sec = 1210/3 =~ 403.3 ft/sec
83 mi/hr =~ 121.7 ft/sec
------------
accel = (403.3 - 121.7)/3 =~ 93.867 ft/sec/sec
----
Some might put a minus sign for slowing down, but that's a matter of reference frames.